# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 6.2. Structure constants under flops
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\to\Delta$ be two quasi-projective open Kulikov degenerations differing by an M1 flop. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, we have the equality
$N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{h(\beta)}(\mathcal{X}^{\prime})\,,$
where $h$ is the isomorphism given in Equation (6.2).
@@
It follows from the properness of moduli spaces given by Theorem 5.2 that the results proved in [39, §8] on punctured Gromov–Witten invariants of Type III degenerations of K3 surfaces also hold by the same arguments for the punctured Gromov–Witten $N_{pqr}^{\beta}(\mathcal{X})$ of quasi-projective open Kulikov degenerations. By [39, Theorem 8.15], the punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be decomposed as a product of punctured Gromov–Witten invariants of the irreducible components $(X^{i},\partial X^{i})$ of $\mathcal{X}_{0}$ and of toric surfaces (appearing as irreducible components of a large enough base change of $\mathcal{X}\rightarrow\Delta$). On the other hand, the M1 flop $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ induces an isomorphism between $\mathcal{X}_{0}$ and $\mathcal{X}_{0}^{\prime}$, except in restriction to the irreducible component $X^{v}$ containing the flopping curve $C$, and to the only other irreducible components $X^{v^{\prime}}$ intersecting $C$. Hence, Theorem 6.2 reduces to a local statement comparing punctured Gromov–Witten invariants of $X^{v}\cup X^{v^{\prime}}$ under the blow-down of $C$ in $X_{v}$, and the blow-up of the point $C\cap X^{v^{\prime}}$ in $X^{v^{\prime}}$. Finally, this statement follows from [39, Theorem 8.16] describing the change of the punctured Gromov–Witten invariants of a log Calabi–Yau surfaces under internal blow-up. We refer to [43] for more details. ∎
@@
Let $\mathcal{X}^{\prime}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration obtained from a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$ by an M1 flop along a curve $C$ in $\mathcal{X}$. The curve $C$ corresponds to an edge $e_{C}$ of $\mathscr{P}$, which is a diagonal in a quadrilateral $Q$. The polyhedral decomposition $\mathscr{P}^{\prime}$ of $P$ defined by $\mathcal{X}^{\prime}$ is obtained from $\mathscr{P}$ by flip of $e_{C}$ in $Q$, that is, by removing the diagonal $e_{C}$ of $Q$ and introducing the opposite diagonal $e_{C^{\prime}}$ of $Q$. Let $c_{Q}\in C(\mathscr{P})_{\mathbb{Z}}$ be the integral primitive generator of the ray in $C(\mathscr{P})$ passing through the intersection point of $e_{C}$ and $e_{C^{\prime}}$ in $P\subset C(P)$. Geometrically, $c_{Q}$ is the divisorial valuation defined by the exceptional divisor $E$ of the blow-up $\mathcal{Z}$ of the ordinary double point obtained by contracting $C$. Denote by $\widetilde{C(\mathscr{P})}$ the dual intersection complex of $\mathcal{Z}$, that is, the cone complex obtained from $C(\mathscr{P})$ by adding the cone over $e_{C^{\prime}}$. Let
@@
As $\mathcal{Z}$ is a log modification of $\mathcal{X}$, induced by the subdivision $\widetilde{C(\mathscr{P})}$ of $C(\mathscr{P})$, we have by [55, §10]
@@
In fact, arguing as in the proof of [39, Lemma 8.7], there exists a unique $\widetilde{\beta}$ such that $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})\neq 0$, and so we have $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})$. Similarly, we have $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})=N_{pqr}^{p^{\prime}_{\star}\widetilde{\beta}}(\mathcal{X}^{\prime})$. Hence, in order to prove Theorem 6.4, it is enough to show that $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+m_{pqr}C$. This formula holds since, by geometry of the flop, we have $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+(E\cdot\widetilde{\beta})C$, and we have $E\cdot\widetilde{\beta}=\varphi_{E}(p)+\varphi_{E}(q)-\varphi_{E}(r)$ by Lemma 6.7. ∎
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. As in [44, Definition 1.8], let $\varphi$ be the $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-valued multivalued PL function on $C(P)$ with kink $[C]\in NE(\mathcal{X}/\overline{\mathcal{X}})$ across the codimension one face of $C(\mathscr{P})$ corresponding to the irreducible component $[C]$ of the double locus of $\mathcal{X}_{0}$. By [44, Construction 1.14], $\varphi$ determines a cone complex $(C(P)_{\varphi},C(\mathscr{P})_{\varphi})$, and a map of cone complexes $\pi:(C(P)_{\varphi},C(\mathscr{P})_{\varphi})\rightarrow(C(P),C(\mathscr{P}))$ such that $\pi:C(P)_{\varphi}\rightarrow C(P)$ is a $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-torsor, and $\varphi$ can naturally be viewed as a section of $\pi$.
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\rightarrow\Delta$ be two quasi-projective open Kulikov degenerations with $\overline{\mathcal{X}}=\overline{\mathcal{X}}^{\prime}$, and $\varphi$ and $\varphi^{\prime}$ the corresponding multivalued PL functions. It follows from [49, §2.1] that we have a natural homeomorphism $C(P)_{\varphi}\simeq C(P^{\prime})_{\varphi^{\prime}}$, preserving the integral points, but not the cone decompositions $C(\mathscr{P})_{\varphi}$ and $C(\mathscr{P}^{\prime})_{\varphi^{\prime}}$ in general. In particular, the difference $\varphi^{\prime}-\varphi$ can be viewed as an ordinary $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-valued function on $C(P)_{\mathbb{Z}}\simeq C(P^{\prime})_{\mathbb{Z}}$.
@@ 6.3. The finiteness property
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. An effective divisor $F$ on $\mathcal{X}$ is said to be good if $F$ is supported on $\mathcal{X}_{0}$ and satisfies the following conditions:
@@
    $-F$ is nef, and for any compact irreducible curve $C\subset\mathcal{X}_{0}$ we have $-F\cdot C\geq 1$, except for finitely many disjoint curves $C$ which are either internal exceptional curves in an irreducible component of $\mathcal{X}_{0}$, or irreducible components $D_{ij}=X^{i}\cap X^{j}$ of the double locus of $\mathcal{X}_{0}$, satisfying $D_{ij}|_{X^{i}}^{2}=D_{ij}|_{X^{j}}^{2}=-1$,
@@
If $x$ is contained in the interior of a 2-dimensional face of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. Moreover, all the walls passing through $x$ are trivial modulo $J$ by Lemma 6.9, and so the walls non-trivial modulo $J^{k}$ are uniquely determined by consistency from the walls non-trivial modulo $J^{k-1}$, and so the finiteness and polynomiality of the walls non-trivial modulo $J^{k}$ follows from the induction hypothesis and [38, Theorem 6.38].
If $x$ is contained in the interior of an edge of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is still given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. The only difference with the previous case is that some walls passing through $x$ might be trivial modulo $J$. However, these walls are of the explicit form given by Lemma 6.9, and so the result still holds by [38, Proposition 6.47].
It remains to consider the case where $x=v$ is a vertex of $\mathscr{P}$, corresponding to an irreducible component $X^{v}$ of $\mathcal{X}_{0}$. It follows from [39, Theorem 8.15] that the punctured Gromov–Witten invariants contributing to the walls $(W,f_{W})$ with initial point $x=v$ and non-trivial modulo $J^{k}$ can be expressed explicitly in terms of punctured Gromov–Witten invariants $N_{\tau}^{\beta}$ of $(X^{v},\partial X_{v})$ and of the walls non-trivial modulo $J^{k-1}$ passing through $v$. More precisely, it follows from [39, Lemma 8.9] that, up to applying corner blow-ups to $(X^{v},\partial X^{v})$, the invariants $N_{\tau}^{\beta}$ are counts of curves $C$ in $X^{v}$ intersecting $\partial X^{v}$ at fixed smooth points $x_{1},\dots,x_{q}$, with multiplicities $m_{1},\dots,m_{q}$, and at a smooth non-fixed point $x_{\mathrm{out}}$ with multiplicity $m_{\mathrm{out}}$. If $q\neq 1$, then, as in the proof of [45, Lemma 4.1], given the points $x_{1}$,…, $x_{q}$, there are only finitely many possibilities for the point $x_{\mathrm{out}}$, depending in a non-trivial way on $x_{1}$,…, $x_{q}$. In particular, for generic $x_{1},\dots,x_{q}$, none of the intersection points of $C$ with $\partial X^{v}$ are intersection of an internal exceptional curve with $\partial X^{v}$, and so $C$ does not contain any internal exceptional curve. If $q=0$ and $x_{\mathrm{out}}$ is an intersection point of an internal exceptional curve $E$ with $\partial X^{v}$, then $\mathcal{O}(C-kE)$ has trivial restriction to $\partial X^{v}$. Since the invariant $N_{\tau}^{\beta}$ is invariant under deformation of $(X^{v},\partial X^{v})$, it follows from [41, Proposition 4.1] that $C=kE$ if $N_{\tau}^{\beta}\neq 0$.
@@
Let $\mathscr{P}^{\prime}$ be the coarsening of $\mathscr{P}$ obtained by removing the edges $e$ such that the corresponding double curve $X^{e}$ of $\mathcal{X}_{0}$ is contained in $G$. Such curves are distinct, and so $\mathscr{P}^{\prime}$ is just obtained by removing diagonals of some quadrilaterals of $\mathscr{P}$. By Equation 5.7, each time a jagged path crosses an edge of $\mathscr{P}^{\prime}$, the corresponding kink is non-trivial modulo $J$, and so the intersection with $-F$ of the curve class attached to the jagged path increases by $1$. Since $-F\cdot\beta=k$ is fixed by $p,q,r$, one deduces that the number of times a jagged path contributing to $N_{pqr}^{\beta}(\mathcal{X})$ crosses an edge of $\mathscr{P}^{\prime}$ is uniformly bounded. Moreover, if $(W,f_{W})$ is a wall generically contained in a 2-dimensional face of $\mathscr{P}^{\prime}$, then $f_{W}=1\mod J$ by Lemma 6.9, and so, by Equation (5.7), the number of times a jagged path crosses such a wall is also uniformly bounded. Hence, we have in fact a uniform bound on the number of times a jagged path can cross a wall. In addition, by Lemma 6.9, there are only finitely many walls non-trivial modulo $J^{k}$, and they all contain only finitely many terms modulo $J^{k}$. Since the jagged paths have fixed asymptotic monomials $z^{p}$ or $z^{q}$, we obtain at every wall-crossing finitely many monomials, and since the number of wall-crossings is uniformly bounded, we conclude that there are finitely many possible jagged paths contributing to structures constants of the form $N_{pqr}^{\beta}(\mathcal{X})$ with fixed $p,q,r$, and so in particular finitely $\beta$ with $N_{pqr}^{\beta}(\mathcal{X})\neq 0$. ∎
@@
    $\mathcal{X}$ has $\mathbb{Q}$-factorial terminal singularities,
@@
    $f:=g\circ h:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is crepant,
@@
    $g:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one,
@@
First recall that by general MMP, [23, Corollary 1.4.3] – see also [58, Corollary 2.6], there exists a crepant $\mathbb{Q}$-factorial terminalization of $\overline{\mathcal{X}}$. Moreover, two such crepant $\mathbb{Q}$-factorial terminalizations are isomorphic in codimension one – see e.g. [60], and so have the same number $e(\overline{\mathcal{X}})$ of exceptional divisors.
@@
Let $\pi\colon\mathcal{X}\to\Delta$ be a quasi-projective open Kulikov degeneration. Then, there exists a quasi-projective open Kulikov degeneration $\pi^{\prime}\colon\mathcal{X}^{\prime}\rightarrow\Delta$, with affinization $\overline{\mathcal{X}}$, such that there exists a good effective divisor $F^{\prime}$ on $\mathcal{X}^{\prime}$.
@@
The result follows immediately from Lemma 6.12, and from the fact that, if $h:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one, then, the exceptional locus of $g$ is a configuration of curves as in Definition 6.8 i). ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. For every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, we have
$h(r)=h(p)+h(q)\,.$

### 30 Mar 2024 v2

@@ 6.2. Structure constants under flops
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\to\Delta$ be two quasi-projective open Kulikov degenerations differing by an M1 flop. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, we have the equality $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{h(\beta)}(\mathcal{X}^{\prime})$, where $h$ is the isomorphism given in Equation (6.2).
@@
It follows from the properness of moduli spaces given by Theorem 5.2 that the results proved in [40, §8] on punctured Gromov–Witten invariants of Type III degenerations of K3 surfaces also hold by the same arguments for the punctured Gromov–Witten $N_{pqr}^{\beta}(\mathcal{X})$ of quasi-projective open Kulikov degenerations. By [40, Theorem 8.15], the punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be decomposed as a product of punctured Gromov–Witten invariants of the irreducible components $(X^{i},\partial X^{i})$ of $\mathcal{X}_{0}$ and of toric surfaces (appearing as irreducible components of a large enough base change of $\mathcal{X}\rightarrow\Delta$). On the other hand, the M1 flop $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ induces an isomorphism between $\mathcal{X}_{0}$ and $\mathcal{X}_{0}^{\prime}$, except in restriction to the irreducible component $X^{v}$ containing the flopping curve $C$, and to the only other irreducible components $X^{v^{\prime}}$ intersecting $C$. Hence, Theorem 6.2 reduces to a local statement comparing punctured Gromov–Witten invariants of $X^{v}\cup X^{v^{\prime}}$ under the blow-down of $C$ in $X_{v}$, and the blow-up of the point $C\cap X^{v^{\prime}}$ in $X^{v^{\prime}}$. Finally, this statement follows from [40, Theorem 8.16] describing the change of the punctured Gromov–Witten invariants of a log Calabi–Yau surfaces under internal blow-up. We refer to [44] for more details. ∎
@@
Let $\mathcal{X}^{\prime}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration obtained from a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$ by an M2 flop along a curve $C$ in $\mathcal{X}$. The curve $C$ corresponds to an edge $e_{C}$ of $\mathscr{P}$, which is a diagonal in a quadrilateral $Q$. The polyhedral decomposition $\mathscr{P}^{\prime}$ of $P$ defined by $\mathcal{X}^{\prime}$ is obtained from $\mathscr{P}$ by flip of $e_{C}$ in $Q$, that is, by removing the diagonal $e_{C}$ of $Q$ and introducing the opposite diagonal $e_{C^{\prime}}$ of $Q$. Let $c_{Q}\in C(\mathscr{P})_{\mathbb{Z}}$ be the integral primitive generator of the ray in $C(\mathscr{P})$ passing through the intersection point of $e_{C}$ and $e_{C^{\prime}}$ in $P\subset C(P)$. Geometrically, $c_{Q}$ is the divisorial valuation defined by the exceptional divisor $E$ of the blow-up $\mathcal{Z}$ of the ordinary double point obtained by contracting $C$. Denote by $\widetilde{C(\mathscr{P})}$ the dual intersection complex of $\mathcal{Z}$, that is, the cone complex obtained from $C(\mathscr{P})$ by adding the cone over $e_{C^{\prime}}$. Let
@@
As $\mathcal{Z}$ is a log modification of $\mathcal{X}$, induced by the subdivision $\widetilde{C(\mathscr{P})}$ of $C(\mathscr{P})$, we have by [56, §10]
@@
In fact, arguing as in the proof of [40, Lemma 8.7], there exists a unique $\widetilde{\beta}$ such that $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})\neq 0$, and so we have $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})$. Similarly, we have $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})=N_{pqr}^{p^{\prime}_{\star}\widetilde{\beta}}(\mathcal{X}^{\prime})$. Hence, in order to prove Theorem 6.4, it is enough to show that $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+m_{pqr}C$. This formula holds since, by geometry of the flop, we have $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+(E\cdot\widetilde{\beta})C$, and we have $E\cdot\widetilde{\beta}=\varphi_{E}(p)+\varphi_{E}(q)-\varphi_{E}(r)$ by Lemma 6.7. ∎
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. As in [45, Definition 1.8], let $\varphi$ be the $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-valued multivalued PL function on $C(P)$ with kink $[C]\in NE(\mathcal{X}/\overline{\mathcal{X}})$ across the codimension one face of $C(\mathscr{P})$ corresponding to the irreducible component $[C]$ of the double locus of $\mathcal{X}_{0}$. By [45, Construction 1.14], $\varphi$ determines a cone complex $(C(P)_{\varphi},C(\mathscr{P})_{\varphi})$, and a map of cone complexes $\pi:(C(P)_{\varphi},C(\mathscr{P})_{\varphi})\rightarrow(C(P),C(\mathscr{P}))$ such that $\pi:C(P)_{\varphi}\rightarrow C(P)$ is a $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-torsor, and $\varphi$ can naturally be viewed as a section of $\pi$.
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\rightarrow\Delta$ be two quasi-projective open Kulikov degenerations with $\overline{\mathcal{X}}=\overline{\mathcal{X}}^{\prime}$, and $\varphi$ and $\varphi^{\prime}$ the corresponding multivalued PL functions. It follows from [50, §2.1] that we have a natural homeomorphism $C(P)_{\varphi}\simeq C(P^{\prime})_{\varphi^{\prime}}$, preserving the integral points, but not the cone decompositions $C(\mathscr{P})_{\varphi}$ and $C(\mathscr{P}^{\prime})_{\varphi^{\prime}}$ in general. In particular, the difference $\varphi^{\prime}-\varphi$ can be viewed as an ordinary $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-valued function on $C(P)_{\mathbb{Z}}\simeq C(P^{\prime})_{\mathbb{Z}}$.
@@ 6.3. The finiteness property
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. An effective divisor $F$ on $\mathcal{X}$ is said to be good if $F$ is supported on $\mathcal{X}_{0}$ and satisfies the following conditions.
@@
    The divisor $-F$ is nef, and for any compact irreducible curve $C\subset\mathcal{X}_{0}$ we have $-F\cdot C\geq 1$, except for finitely many disjoint curves $C$ which are either internal exceptional curves in an irreducible component of $\mathcal{X}_{0}$, or irreducible components $D_{ij}=X^{i}\cap X^{j}$ of the double locus of $\mathcal{X}_{0}$, satisfying $D_{ij}|_{X^{i}}^{2}=D_{ij}|_{X^{j}}^{2}=-1$.
@@
If $x$ is contained in the interior of a 2-dimensional face of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. Moreover, all the walls passing through $x$ are trivial modulo $J$ by Lemma 6.9, and so the walls non-trivial modulo $J^{k}$ are uniquely determined by consistency from the walls non-trivial modulo $J^{k-1}$, and so the finiteness and polynomiality of the walls non-trivial modulo $J^{k}$ follows from the induction hypothesis and [39, Theorem 6.38].
If $x$ is contained in the interior of an edge of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is still given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. The only difference with the previous case is that some walls passing through $x$ might be trivial modulo $J$. However, these walls are of the explicit form given by Lemma 6.9, and so the result still holds by [39, Proposition 6.47].
It remains to consider the case where $x=v$ is a vertex of $\mathscr{P}$, corresponding to an irreducible component $X^{v}$ of $\mathcal{X}_{0}$. It follows from [40, Theorem 8.15] that the punctured Gromov–Witten invariants contributing to the walls $(W,f_{W})$ with initial point $x=v$ and non-trivial modulo $J^{k}$ can be expressed explicitly in terms of punctured Gromov–Witten invariants $N_{\tau}^{\beta}$ of $(X^{v},\partial X_{v})$ and of the walls non-trivial modulo $J^{k-1}$ passing through $v$. More precisely, it follows from [40, Lemma 8.9] that, up to applying corner blow-ups to $(X^{v},\partial X^{v})$, the invariants $N_{\tau}^{\beta}$ are counts of curves $C$ in $X^{v}$ intersecting $\partial X^{v}$ at fixed smooth points $x_{1},\dots,x_{q}$, with multiplicities $m_{1},\dots,m_{q}$, and at a smooth non-fixed point $x_{\mathrm{out}}$ with multiplicity $m_{\mathrm{out}}$. If $q\neq 1$, then, as in the proof of [46, Lemma 4.1], given the points $x_{1}$,…, $x_{q}$, there are only finitely many possibilities for the point $x_{\mathrm{out}}$, depending in a non-trivial way on $x_{1}$,…, $x_{q}$. In particular, for generic $x_{1},\dots,x_{q}$, none of the intersection points of $C$ with $\partial X^{v}$ are intersection of an internal exceptional curve with $\partial X^{v}$, and so $C$ does not contain any internal exceptional curve. If $q=0$ and $x_{\mathrm{out}}$ is an intersection point of an internal exceptional curve $E$ with $\partial X^{v}$, then $\mathcal{O}(C-kE)$ has trivial restriction to $\partial X^{v}$. Since the invariant $N_{\tau}^{\beta}$ is invariant under deformation of $(X^{v},\partial X^{v})$, it follows from [42, Proposition 4.1] that $C=kE$ if $N_{\tau}^{\beta}\neq 0$.
@@
Let $\mathscr{P}^{\prime}$ be the coarsening of $\mathscr{P}$ obtained by removing the edges $e$ such that the corresponding double curve $X^{e}$ of $\mathcal{X}_{0}$ is contained in $G$. Such curves are distinct, and so $\mathscr{P}^{\prime}$ is just obtained by removing diagonals of some quadrilaterals of $\mathscr{P}$. By Equation (5.7), each time a jagged path crosses an edge of $\mathscr{P}^{\prime}$, the corresponding kink is non-trivial modulo $J$, and so the intersection with $-F$ of the curve class attached to the jagged path increases by $1$. Since $-F\cdot\beta=k$ is fixed by $p,q,r$, one deduces that the number of times a jagged path contributing to $N_{pqr}^{\beta}(\mathcal{X})$ crosses an edge of $\mathscr{P}^{\prime}$ is uniformly bounded. Moreover, if $(W,f_{W})$ is a wall generically contained in a 2-dimensional face of $\mathscr{P}^{\prime}$, then $f_{W}=1\mod J$ by Lemma 6.9, and so, by Equation (5.7), the number of times a jagged path crosses such a wall is also uniformly bounded. Hence, we have in fact a uniform bound on the number of times a jagged path can cross a wall. In addition, by Lemma 6.9, there are only finitely many walls non-trivial modulo $J^{k}$, and they all contain only finitely many terms modulo $J^{k}$. Since the jagged paths have fixed asymptotic monomials $z^{p}$ or $z^{q}$, we obtain at every wall-crossing finitely many monomials, and since the number of wall-crossings is uniformly bounded, we conclude that there are finitely many possible jagged paths contributing to structures constants of the form $N_{pqr}^{\beta}(\mathcal{X})$ with fixed $p,q,r$, and so in particular finitely $\beta$ with $N_{pqr}^{\beta}(\mathcal{X})\neq 0$. ∎
@@
    $\mathcal{X}$ has $\mathbb{Q}$-factorial terminal singularities.
@@
    $f:=g\circ h:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is crepant.
@@
    $g:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one.
@@
First recall that by general MMP, [23, Corollary 1.4.3] – see also [59, Corollary 2.6], there exists a crepant $\mathbb{Q}$-factorial terminalization of $\overline{\mathcal{X}}$. Moreover, two such crepant $\mathbb{Q}$-factorial terminalizations are isomorphic in codimension one – see e.g. [61], and so have the same number $e(\overline{\mathcal{X}})$ of exceptional divisors.
@@
Let $\pi\colon\mathcal{X}\to\Delta$ be a quasi-projective open Kulikov degeneration. Then, there exists a quasi-projective open Kulikov degeneration $\pi^{\prime}\colon\mathcal{X}^{\prime}\rightarrow\Delta$, with affinization $\overline{\mathcal{X}}$ and having a good effective divisor.
@@
The result follows immediately from Lemma 6.12, and from the fact that, if $h:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one, then the exceptional locus of $g$ is a configuration of curves as in Definition 6.8 i). ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. For every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, we have $h(r)=h(p)+h(q)$.